OpenStudy (tmagloire1):

If 3x^2 + y^2 = 7 then evaluate the second derivative of y with respect to x when x = 1 and y = 2. Round your answer to 2 decimal places.

OpenStudy (tmagloire1):

@jim_thompson5910

OpenStudy (coconutjj):

In order to evaluate the second derivative, you must first know the first derivative. Do you know how to differentiate implicitly

OpenStudy (tmagloire1):

Yes the first derivative is -3x/y

OpenStudy (coconutjj):

dy/dx = -3x/y right? Now what is d/dx(-3x/y)

OpenStudy (coconutjj):

\[\frac{ -3y-(-3x)\frac{ dy }{ dx } }{ y^2 }\]

OpenStudy (tmagloire1):

oh i tustve messed up somewhere but ok i understand that part

OpenStudy (coconutjj):

now what does dy/dx equal to ?

OpenStudy (tmagloire1):

what do you mean? when i plug in 1 and 2?

OpenStudy (coconutjj):

We just solved for dy/dx, dy/dx = -3x/y remember ?

OpenStudy (tmagloire1):

yeah

OpenStudy (coconutjj):

so plug it in and simplify

OpenStudy (tmagloire1):

plug it into the first or second derivative?

OpenStudy (coconutjj):

well, which one seems more logical? 1. Plugging in to 1st derivative solving for second derivative 2. Plugging into 2nd derivative solving for second derivative

OpenStudy (tmagloire1):

I see the 1st

OpenStudy (tmagloire1):

-3(1)/(2) = -3/2

OpenStudy (coconutjj):

No... not quite.. I don't know how you would approach the 1st. 2nd one is correct

OpenStudy (tmagloire1):

-3(2)+3(1)/2^2 = -6+3/4 = -3/4

OpenStudy (coconutjj):

Remember it is \[\frac{ d^2y }{dx^2 }\] = ...

OpenStudy (coconutjj):

\[\frac{ -3y - (3x)(\frac{ -3x }{ y }) }{ y^2 }\]

OpenStudy (tmagloire1):

-3(2)+3(1)(-3/2)/2^2 = -21/8

OpenStudy (coconutjj):

you should be correct if you did the calculations correct.

OpenStudy (tmagloire1):

ok thank you for your help!